# Group Relative Attention Guidance for Image Editing

### v1

@@ B.1 Kontext
We provide additional kontext model feature distribution statistics corresponding to Figures 2, 4, and 5 in the main paper. Consistent with the experiments presented in the main text, we analyze different image editing samples (IDs) across various denoising steps and model layers to examine the correlation between feature distributions and these three factors. Figure S1 presents direct visualizations of the feature distributions, where the TokenNumber dimension is downsampled by a factor of 4 and the Dim dimension by a factor of 2. Figure S2 shows aggregating different tokens along the sequence dimension. Figures S3S6 illustrate the mean and variance of token-wise feature distributions across different attention heads, corresponding to the different embedding features.
@@
C The Use of Large Language Models
In this work, we use Gemini-2.5-pro to aid in the writing process. Specifically, the model was used to improve the grammatical structure, refine sentence phrasing, and enhance the overall readability of the text. The core scientific content, methodologies, and conclusions presented in this paper are the original work of the authors. The use of the LLM was restricted to a tool for language enhancement.

### v2

@@ Appendix E Additional Feature Visualization
We provide additional Kontext and Qwen-Edit model feature distribution statistics corresponding to Figures 2, 4, and 5 in the main paper. Consistent with the experiments presented in the main text, we analyze different image editing samples (IDs) across various denoising steps and model layers to examine the correlation between feature distributions and these three factors. Figure S9, Figure S15 presents direct visualizations of the feature distributions, where the TokenNumber dimension is downsampled by a factor of 4 and the Dim dimension by a factor of 2. Figure S10, Figure S16 shows aggregating different tokens along the sequence dimension. Figures S11S14, Figures S17S20 illustrate the mean and variance of token-wise feature distributions across different attention heads, corresponding to the different embedding features.
E.1 Kontext Embedding Visualization
@@
E.2 Qwen-Edit Embedding Visualization